# Vortices and Factorization

### v2

@@ 5 Darboux Transformation and Equilibrium Configurations
As we mentioned above, sequences of the equilibrium configurations can be produced with help of Darboux transformations. To construct them, we need to introduce the factorizable function corresponding to a static configuration (6)
$\psi(z)=C\prod_{i=1}^{N}(z-z_{i})^{{\cal Q}_{i}},$ (46)
where $C$ is a non-zero constant and $Q_{i}$ are not necessarily integers.
Now, we will show that if the Darboux transform $\hat{\psi}$ of $\psi$ has a similar form, i.e.
$\hat{\psi}(z)=\hat{C}\prod_{i=1}^{\hat{N}}(z-\hat{z}_{i})^{\hat{\cal Q}_{i}},$ (47)
then $\hat{\psi}$ also corresponds to a static configuration. For this we introduce the Schrodinger operator $H$ whose kernel contains $\psi$, i.e.
$H\psi=(-\partial_{z}^{2}+u)\psi=0.$
Then, using (46), we obtain
$u=\frac{\psi^{\prime\prime}}{\psi}=(\log\psi)^{\prime\prime}+((\log\psi)^{\prime})^{2}=-\sum_{i}\frac{{\cal Q}_{i}}{(z-z_{i})^{2}}+\sum_{i,j}\frac{{\cal Q}_{i}{\cal Q}_{j}}{(z-z_{i})(z-z_{j})}.$
Applying the partial fraction decompositions in the last sum we get
$u=\sum_{i}\frac{{\cal Q}_{i}({\cal Q}_{i}-1)}{(z-z_{i})^{2}}+2\sum_{i}\frac{{\cal Q}_{i}}{(z-z_{i})}\sum_{j\not=i}\frac{{\cal Q}_{j}}{(z_{i}-z_{j})}.$
Due to the equilibrium condition (6) the last term in the above equation vanishes and potential is a sum of the second order poles
$u=\sum_{i}\frac{{\cal Q}_{i}({\cal Q}_{i}-1)}{(z-z_{i})^{2}},$ (48)
i.e. the equilibrium condition is equivalent to absence of the first-order poles in the potential.
Now, to prove that the Darboux transform $\hat{\psi}$ of $\psi$ of the form (47) also correspond to an equilibrium configuration, i.e.
$\sum_{j=1,j\not=i}^{\hat{N}}\frac{\hat{\cal Q}_{j}}{\hat{z}_{i}-\hat{z}_{j}}=0,\quad i=1\dots\hat{N},$
we have to show that potential $\hat{u}$ of the Darboux transformed Schrodinger operator $\hat{H}$ is also a sum of the second order poles. Indeed, the Darboux transformed potential equals (see (33))
$\hat{u}=u-2(\log\psi)^{\prime\prime},$
and from (46) and (48) we obtain
@@
i.e. $\hat{u}$ is a sum of second order poles and therefore $\hat{\psi}$ of the form (47) corresponds to a static configuration.
@@
from (49), we see that the Darboux transformation acts on charges either of two ways
@@
i.e. it either (i) inverts a charge or (ii) increments a charge by 1. In particular, it can create new charges with ${\cal Q}_{i}=1$ and annihilate charges with ${\cal Q}_{i}=-1$. In the case of the Adler-Moser polynomials in generic configurations, it annihilates negative charges, inverts positive charges, and creates positive charges at new positions.
6 Configurations Related to Bi-spectral Systems
In section 3 we showed that the static configurations of the system consisting of charges of values ±$\pm 1$ (i.e. when $\Lambda=1$) correspond to zeros of the Adler-Moser polynomials $P_{n}$ and $P_{n+1}$. In section 4 these polynomials were obtained with the help of sequence of the Darboux transformations (35) at zero energy level. Corresponding sequence of $H_{i}$ is started from the free Schrodinger operator $H_{0}=-\partial_{z}^{2}$. This sequence is known to constitute the “odd” family of bi-spectral operators [^7] The differential operators $H$ in $z$ is bi-spectral if there exists a differential operator $B$ in $k$ and a common eigenfunction, such that $H\psi(z,k)=f(k)\psi(z,k)$, $B\psi(z,k)=g(z)\psi(z,k).$ For the even family of Schrodinger operators, the common eigenspace of $H$ and $B$ is two-dimensional. It is one-dimensional in the odd case. [17]. There exists another sequence of the Darboux transforms consisting of “even” family, of bi-spectral operators [17]. It starts with
@@ 6 Configurations Related to Bi-spectral Systems
In both “even” and “odd” cases we present elements of the sequence of Darboux-transforms of kernels of $H_{i}$
@@
According to section 4 (see eq. (33)) the Darboux transform $\psi_{i+1}$ of $\psi_{i}$ is
@@
From this and (52) we get the differential recurrent relation for $P_{i}$
@@
In the “odd”, i.e. the Adler-Moser case $P_{0}=1$, $P_{1}=z$ (see eq. (22)). In the “even” case $H_{0}[z^{1/2}]=0$ (i.e. $\psi_{0}=z^{1/2}$ see (51)) and, according to (52)
@@
Note that, for convenience, we set the normalization constants $C_{i}$ in (54) to values at which $P_{i}(z)$ are monic [^8] In the even case, this differs from the original normalization of [17], where same $C_{i}=2i+1$ were chosen for both families.. For “even” family, from (54,55) we get $P_{2}=z^{2}+t_{1}$. So, using (52), we obtain the sequence of length 2 depending on one parameter $t_{1}$:
$\psi_{0}=z^{1/2},\quad\psi_{1}=\frac{z^{2}+t_{1}}{z^{1/2}}.$ (56)
Another application of the recurrence relation (54) with $t_{1}\not=0$ would produce a logarithmic term, so to obtain $\psi_{2}$ and $\psi_{3}$ of the form (46) we restrict $t_{1}$ to be zero and pick a new integration constant $t_{2}$:
$P_{3}=z^{9/2}+t_{2}z^{1/2},\quad P_{4}=z^{8}+6t_{2}z^{4}+t_{3}z^{2}-3t_{2}^{2}.$ (57)
Now, we get the sequence of length 4:
$\psi_{0}=z^{1/2},\quad\psi_{1}=z^{3/2},\quad\psi_{2}=\frac{z^{4}+t_{2}}{z^{3/2}},\quad\psi_{3}=\frac{z^{8}+6t_{2}z^{4}+t_{3}z^{2}-3t_{2}^{2}}{z^{9/2}+t_{2}z^{1/2}}$
depending on two parameters $t_{2}$,$t_{3}$.
Next, to eliminate logarithmic terms in $P_{5}$ and $P_{6}$, one has to set both $t_{1}=t_{2}=0$ and we get a sequence $\psi_{0},\psi_{1},\psi_{2},\psi_{3},\psi_{4},\psi_{5}$ (of length 6) depending on three parameters $t_{3}$, $t_{4}$ and $t_{5}$. And so forth.
Sequences of the even family have finite lengths (i.e. they terminate) due to the presence of charge of the third specie at the origin: At the beginning of sequence, charge $1/2$ is placed at $z=0$. Then, according to (50), the following transformations are produced with this charge: It is incremented by $1$ during each of the $m-1$ first Darboux transformations. It is inverted by the $m$th transformation and becomes equal to $-m+1/2$. Then, again, it is incremented during the next $m-1$ steps, until it reaches value $-1/2$. This charge would create a logarithmic singularity in (53) at the next step, so sequence terminates.
Closing this section we note that similarly to the odd family, the wronskian representation of $\psi_{i}$ exists also for the even family. It will be given in section 11 (eqs. (88) and (89)).
@@ 7 KWCC Transformation and Terminating Configurations
Let us now return to the sequence of transformations of rational functions (whose primitives are also rational functions) corresponding to  [^9] Or corresponding to the chain of solutions $\dots\to p_{i-1},q_{i}\to p_{i},q_{i}\to p_{i},q_{i+1}\to p_{i+1},q_{i+1}\to\dots$ of (8). iterations (18,19)
@@
Denoting function $\frac{q_{i}^{2\Lambda}}{p_{i-1}^{2}}$ from the above chain as $\psi^{2}$ and the next function $\frac{p_{i}^{2/\Lambda}}{q_{i}^{2}}$ as $\tilde{\psi}^{2}$, from (19) we get the relationships between them
@@
Next transformation in the sequence (58) relates $\psi^{2}=\frac{p_{i}^{2/\Lambda}}{q_{i}^{2}}$ and $\tilde{\psi}^{2}=\frac{q_{i+1}^{2\Lambda}}{p_{i}^{2}}$. Using (18) we also arrive to transformation (59), this time with $\gamma=\Lambda$.
Therefore, the chain (58) can be constructed by the iterative applications of following transformations
@@
    Exponentiation $\tilde{\psi}=\hat{\psi}^{\gamma}$.
@@
From the above and (50) it follows that KWCC transformation (59) acts on charges either of two ways
$\tilde{\cal Q}=-\gamma{\cal Q},\quad\tilde{\cal Q}=\gamma\left({\cal Q}+1\right).$ (63)
Obviously, the KWCC transformations (60,61,62) generate all classes of configurations considered above, i.e. those related to the odd and even bi-spectral families as well as to the $\Lambda=2$ case.
@@
where $\tau_{i}$ is a polynomial of an integer or half-integer degree $d_{i}$. Here, “polynomial of half-integer degree” in $z$ stands for a polynomial in $z$ times $z^{1/2}$. Then from (60) and (61) we get
@@
These are actually relations (21), but now $d_{i}$ are not necessarily degrees (15) or (17) of the main sequences. This is because (64, 65) are derived from the KWCC chain, rather than from the bi-linear equation (8). Because of this fact, any polynomial solution of (65) that starts from a static configuration (e.g. from a single vortex) also corresponds to a sequence of static configurations.
Apart from infinite main sequences of the $\Lambda=1$ and $\Lambda=2$ cases, recurrence relations (64,65) allow to find sequences of configurations of finite lengths, i.e. terminating configurations. The “even” bi-spectral family of section 6, i.e. solutions of (64,65) corresponding to $\gamma_{i}=1$, $d_{1}=1/2$, is a family of terminating configurations for $\Lambda=1$.
As an example of family of terminating configurations for $\Lambda=2$, we can take the case $\gamma_{0}=2$, $d_{1}=2$, i.e. $\tau_{1}=z^{2}$. Here are several first terminating sequences of $\tau_{i}$:
$\begin{array}[]{c}1,z^{2},z^{3}+t_{1}\\ 1,z^{2},z^{3},z^{2}(z^{9}+t_{2}),z^{9}+t_{3}z^{3}-2t_{2}\\ 1,z^{2},z^{3},z^{11},z^{3}(z^{6}+t_{3}),z^{2}(z^{24}+\frac{20}{3}z^{18}t_{3}+30z^{12}t_{3}^{2}+z^{9}t_{4}-20z^{6}t_{3}^{3}-\frac{5}{3}t_{3}^{4}),\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad z^{18}+15z^{12}t_{3}+z^{9}t_{5}-45z^{6}t_{3}^{2}+\frac{2t_{3}t_{5}-3t_{4}}{2}z^{3}+5t_{3}^{3}\end{array}$
Similarly to the even bi-spectral family, here, a charge of the third specie undergoes transformations (63) at $z=0$ (this time with $\gamma\in\{1/2,2\}$) until the logarithmic singularity is encountered.
One can compute terminating sequences by solving (65) for $\tau_{i+1}$ as for a monic polynomial with unknown coefficients. Compatibility of resulting system of algebraic equations (which is linear in coefficients of $\tau_{i+1}$) imposes restrictions on free parameters of $\tau_{i}$. For sequences related to the even bi-spectral family there exists a wronskian representation (see section 11).
We recall that the main sequences exhaust all possible no-common root polynomial solutions of (8) as well as the collapse solutions for $\Lambda=1$ and $\Lambda=2$ cases. Solutions of (65) whose degrees are not in the main sequences do not satisfy bi-linear equations (8). They satisfy its generalization.
In more detail: Let us first consider sequence of the Darboux transforms at zero energy level that starts from $H_{0}=-\partial_{z}^{2}+u_{0}$, with non-zero $u_{0}\not=0$. At the $n$th step of the chain we have
$u_{n}=u_{0}-2(\log\tau_{n})^{\prime\prime},\quad\varkappa_{n}=\frac{\tau_{n+1}}{\tau_{n}},\quad\tau_{0}=1,\quad H_{n}\varkappa_{n}=0,$
@@
In fact, this generalization is a particular case of the tri-linear equation for three species of vortices with ${\cal Q}_{i}\in\{-1,1,a\}$, where a single vortex of the third specie of strength $a$ is placed at the origin $z=0$ [15, 33]. In general, for multiple species we have the multi-linear equation (see e.g. [25, 32])
$\sum_{i}\Lambda_{i}^{2}\frac{{\cal P}_{i}^{\prime\prime}}{{\cal P}_{i}}+2\sum_{i<j}\Lambda_{i}\Lambda_{j}\frac{{\cal P}_{i}^{\prime}}{{\cal P}_{i}}\frac{{\cal P}_{j}^{\prime}}{{\cal P}_{j}}=0,$
where sums run over all species, $\Lambda_{i}$ is a strength of the $i$th specie and ${\cal P}_{i}$ is polynomial whose roots are coordinates of vortices of this specie. In the case of configurations related to the even bi-spectral family ${\cal P}_{1}=p$, ${\cal P}_{2}=q$, ${\cal P}_{3}=z$. In similar situation, when ${\cal Q}_{i}\in\{-1,\Lambda,\Lambda a\}$, we have the following generalization of (8)

### v3

@@ 5 Darboux Transformation and Equilibrium Configurations
According to (43), in the case of Adler-Moser polynomials, iterations of Darboux transformations produce Schrodinger operators with potentials $u_{n}=-2(\log P_{n})^{\prime\prime}$ that are free of simple poles. Corresponding zero-level eigenfunctions $\varkappa_{n}=P_{n+1}/P_{n}$ are of the form $\prod_{i}(z-z_{i})^{{\cal Q}_{i}}$, where ${\cal Q}_{i}=\pm 1$ are the values of charges in equilibrium, and $z_{i}$ are their positions.
Let us consider a more general situation, by introducing a factorizable eigenfunction
$\psi(z)=\prod_{i=1}^{N}(z-z_{i})^{{\cal Q}_{i}},$ (47)
where, for the moment, we treat $z_{i}$ and ${\cal Q}_{i}$ as some unspecified numbers. Let us now look for all possible ${\cal Q}_{i}$ and $z_{i}$ for which $\psi$ satisfies the Schrodinger equation with the potential being free of simple poles:
$H\psi=(-\partial_{z}^{2}+u)\psi=0,\quad u\,\,\,\text{is free of simple poles}.$
Since
$\frac{H\psi}{\psi}=u-(\log\psi)^{\prime\prime}-((\log\psi)^{\prime})^{2}=u-\sum_{i}\frac{{\cal Q}_{i}}{(z-z_{i})^{2}}-\sum_{i,j}\frac{{\cal Q}_{i}{\cal Q}_{j}}{(z-z_{i})(z-z_{j})},$ (48)
the Schrodinger equation is satisfied if and only if the expression on the right-hand side of (48) vanishes identically in $z$. Applying the partial-fraction decomposition, we rewrite this expression as
$u-\sum_{i}\frac{{\cal Q}_{i}({\cal Q}_{i}-1)}{(z-z_{i})^{2}}-\sum_{i}\frac{2{\cal Q}_{i}\sum_{j\not=i}\frac{{\cal Q}_{j}}{z_{i}-z_{j}}}{z-z_{i}}.$ (49)
For $u$ that is free of simple poles, (49) can vanish identically in $z$ only if all residues $-2{\cal Q}_{i}\sum_{j\not=i}\frac{{\cal Q}_{j}}{z_{i}-z_{j}}$ equal zero, i.e., only if $\sum_{j\not=i}\frac{{\cal Q}_{j}}{z_{i}-z_{j}}=0$ for all $1\leq i\leq N$, which is precisely equilibrium conditions (6). Hence, if $u$ is free of simple poles, ${\cal Q}_{i}$ and $z_{i}$ must satisfy (6). The converse also follows from (49), so the equilibrium condition is equivalent to the absence of simple poles in the potential. From (49), we also obtain the potential
$u=\sum_{i}\frac{{\cal Q}_{i}({\cal Q}_{i}-1)}{(z-z_{i})^{2}}.$ (50)
Now, suppose that there exists a factorizable Darboux transform $\hat{\psi}$ of $\psi$:
$\hat{\psi}(z)=\prod_{i=1}^{\hat{N}}(z-\hat{z}_{i})^{\hat{\cal Q}_{i}}.$ (51)
Then $\hat{\psi}$ also corresponds to a static configuration.
Here $\hat{N}$ charges $\hat{\cal Q}_{1},\dots,\hat{\cal Q}_{N}$ are in equilibrium at points $\hat{z}_{1},\dots,\hat{z}_{\hat{N}}$. Indeed, equilibrium conditions for $\hat{\cal Q}_{i}$, $\hat{z}_{i}$ are satisfied if the potential $\hat{u}$ of the Darboux transformed Schrodinger operator $\hat{H}$ is free of simple poles. According to (31), transform of $u$ equals
$\hat{u}=u-2(\log\psi)^{\prime\prime}.$ (52)
Then, from (47) and (50) we obtain
@@
i.e. $\hat{u}$ is a sum of second order poles. Therefore, $\hat{\psi}$, of the form (51), corresponds to a static configuration.
@@
from (53), we see that the Darboux transformation can act on charges in two ways
@@
i.e., it either (i) inverts a charge or (ii) increments a charge by 1. In particular, it can create new charges out of “zero charges” $\hat{\cal Q}=0+1=1$ and annihilate charges with ${\cal Q}=-1$, $\hat{\cal Q}=-1+1=0$. In the case of the Adler-Moser polynomials in generic configurations, it annihilates negative charges, inverts positive charges, and creates positive charges at new positions.
We also note that, when all charges belong to the set
${\cal Q}_{i}\in\left\{-1,\frac{1}{2},1,\frac{3}{2},2,\dots\right\},$
[…]
The presence of charges ${\cal Q}_{i}=-1/2$ always leads to logarithmic terms in $\hat{\psi}$.
6 Configurations Related to Even Bispectral Family
In this section, we review a family of factorizable eigenfunctions, generated through Darboux transformations, that correspond to the equilibria of three species of charges [22, 29, 42].
We recall that the sequence $H_{i}$ of Schrodinger operators, corresponding to the Adler-Moser polynomials begins with the free Schrodinger operator $H_{0}=-\partial_{z}^{2}$. This sequence is known to constitute the “odd” family of bi-spectral operators [^14] The differential operator $H$ in $z$ is bi-spectral if there exists a differential operator $B$ in $k$ and a common eigenfunction, such that $H\psi(z,k)=f(k)\psi(z,k)$, $B\psi(z,k)=g(z)\psi(z,k).$ For the even family of Schrodinger operators, the common eigenspace of $H$ and $B$ is two-dimensional, while in the odd case, it is generally one-dimensional. [22]. There also exists another sequence of the Darboux transforms, which forms the “even” family of bi-spectral operators [22]. This sequence starts with the operator
@@ 6 Configurations Related to Even Bispectral Family
In both the “even” and “odd” cases we present elements of the sequence of Darboux-transforms of kernels of $H_{i}$:
@@
According to section 4 (see eq. (34)), the Darboux transform $\psi_{i+1}$ of $\psi_{i}$ is
@@
From this and (56), we derive the differential recurrent relation for $P_{i}$:
@@
In the “odd” (Adler-Moser) case $P_{0}=1$, $P_{1}=z$ (see eq. (23)). In the “even” case $H_{0}[z^{1/2}]=0$ (i.e. $\psi_{0}=z^{1/2}$ see (55)) and, according to (56)
@@
Note that, for convenience, we set the normalization constants $C_{i}$ in (58) to values that make $P_{i}(z)$ monic [^15] In the even case, this differs from the original normalization of [22], where same $C_{i}=2i+1$ were chosen for both families.. For the “even” family, from (58,59) we get $P_{2}=z^{2}+s_{1}$. Thus, using (56), we obtain the sequence of length 2 depending on one parameter $s_{1}$:
$\psi_{0}=z^{1/2},\quad\psi_{1}=\frac{z^{2}+s_{1}}{z^{1/2}}.$ (60)
Another application of the recurrence relation (58) with $s_{1}\not=0$ would produce a logarithmic term. Therefore, to obtain $\psi_{2}$ and $\psi_{3}$ of the form (47), we restrict $s_{1}$ to be zero and pick a new integration constant $s_{2}$:
$P_{3}=z^{9/2}+s_{2}z^{1/2},\quad P_{4}=z^{8}+6s_{2}z^{4}+s_{3}z^{2}-3s_{2}^{2}.$ (61)
Thus, we obtain the sequence of length 4:
$\psi_{0}=z^{1/2},\quad\psi_{1}=z^{3/2},\quad\psi_{2}=\frac{z^{4}+s_{2}}{z^{3/2}},\quad\psi_{3}=\frac{z^{8}+6s_{2}z^{4}+s_{3}z^{2}-3s_{2}^{2}}{z^{9/2}+s_{2}z^{1/2}}$
depending on two parameters, $s_{2}$ and $s_{3}$.
Next, to eliminate logarithmic terms in $P_{5}$ and $P_{6}$, one has to set both $s_{1}=s_{2}=0$. We then obtain a sequence $\psi_{0},\psi_{1},\psi_{2},\psi_{3},\psi_{4},\psi_{5}$ (of length 6) depending on three parameters $s_{3}$, $s_{4}$, and $s_{5}$. And so forth.
Sequences of the even family have finite lengths (i.e., they terminate) due to the presence of charge of the third species at the origin: Initially, charge $1/2$ is placed at $z=0$. According to (54), the following transformations occur with this charge: It is incremented by $1$ during each of the $m-1$ first Darboux transformations. It is inverted by the $m$th transformation and becomes equal to $-m+1/2$. The charge is then incremented again during the next $m-1$ steps, until it reaches value $-1/2$. This charge would create a logarithmic term in (57) at the next step, so the sequence terminates.
Closing this section, we note that similarly to the odd family, the Wronskian representation of $\psi_{i}$ can also be obtained for the even family by applying the procedure (43,44). This will be presented in Section 13 (eqs. (110) and (111)).
@@ 7 KWCC Transformation and Terminating Configurations
Let us now return to the sequence of transformations of rational functions (whose primitives are also rational functions) corresponding to  [^16] Or to the chain of solutions $\dots\to p_{i-1},q_{i}\to p_{i},q_{i}\to p_{i},q_{i+1}\to p_{i+1},q_{i+1}\to\dots$ of (10). iterations (19,20)
@@
Denoting the function $\frac{q_{i}^{2\Lambda}}{p_{i-1}^{2}}$ from the above chain as $\psi^{2}$ and the next function $\frac{p_{i}^{2/\Lambda}}{q_{i}^{2}}$ as $\tilde{\psi}^{2}$, from (20) we get the relationships between them
@@
The next transformation in the sequence (62) relates $\psi^{2}=\frac{p_{i}^{2/\Lambda}}{q_{i}^{2}}$ and $\tilde{\psi}^{2}=\frac{q_{i+1}^{2\Lambda}}{p_{i}^{2}}$. Using (19), we also arrive to transformation (63), this time with $\gamma=\Lambda$.
Therefore, the chain (62) can be constructed by the iterative applications of the following transformations
@@
    Exponentiation [^17] When $\gamma=-1$, the exponentiation $\psi\to 1/\psi$ is a special case of the zero-level Darboux transformation. In this case, we have composition of the Darboux transformation from $H$ to $\hat{H}$ and of another Darboux transformation, from $\hat{H}$ back to $H$. Thus, for $\gamma=-1$, the KWCC transformation is simply a transformation between solutions of the same equation $H\psi=0$. $\tilde{\psi}=\hat{\psi}^{\gamma}$.
@@
From the above and (54), it follows that KWCC transformation (63) can act on charges in two ways
$\tilde{\cal Q}=-\gamma{\cal Q},\quad\tilde{\cal Q}=\gamma\left({\cal Q}+1\right),$ (67)
which includes the creation of new charges out of “zero charges” at new positions, where $\tilde{\cal Q}=\gamma(0+1)=\gamma$. Obviously, the KWCC transformations (64,65,66) generate all classes of configurations considered above, i.e. those related to the odd and even bi-spectral families as well as to the $\Lambda=2$ case.
@@
where $\tau_{i}$ is a polynomial of an integer or half-integer degree $d_{i}$. Here, a “polynomial of half-integer degree” in $z$ means a polynomial in $z$ times $z^{1/2}$. Then from (64) and (65) we get
@@
These are relations (22), but now $d_{i}$ are not necessarily the degrees (16) or (18) of the main sequences. This is because (68, 69) are derived from the KWCC chain, rather than from the bi-linear equation (10). Because of this, any polynomial solution of (69) that starts from a static configuration (e.g. from a single vortex) also corresponds to a sequence of static configurations.
Apart from the infinite main sequences of the $\Lambda=1$ and $\Lambda=2$ cases, recurrence relations (68,69) allow one to find sequences of configurations of finite lengths, i.e., terminating configurations. The “even” bi-spectral family of section 6, i.e., solutions of (68,69) corresponding to $\gamma_{i}=1$, $d_{1}=1/2$, is a family of terminating configurations for $\Lambda=1$. In this case, (69) are relations (58).
In the $\Lambda=2$ case, Eqs. (69) are, modulo normalization factors, relations (25). As an example of a family of terminating configurations for $\Lambda=2$, we can take the case $\gamma_{0}=2$, $d_{1}=2$, i.e., $\tau_{1}=z^{2}$. Here are several first terminating sequences of $\tau_{i}$:
$\begin{array}[]{c}1,z^{2},z^{3}+s_{1}\\ 1,z^{2},z^{3},z^{2}(z^{9}+r_{1}),z^{9}+s_{2}z^{3}-2r_{1}\\ 1,z^{2},z^{3},z^{11},z^{3}(z^{6}+s_{2}),z^{2}(z^{24}+\frac{20}{3}z^{18}s_{2}+30z^{12}s_{2}^{2}+z^{9}r_{2}-20z^{6}s_{2}^{3}-\frac{5}{3}s_{2}^{4}),\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad z^{18}+15z^{12}s_{2}+z^{9}s_{3}-45z^{6}s_{2}^{2}+\frac{2s_{2}s_{3}-3r_{2}}{2}z^{3}+5s_{2}^{3}\end{array}$
Similarly to the even bi-spectral family, here, a charge of the third species undergoes transformations (67) at $z=0$ (this time with $\gamma\in\{1/2,2\}$) until the logarithmic singularity is encountered in (64).
One can compute terminating sequences by solving (69) for $\tau_{i+1}$ as for a monic polynomial with unknown coefficients. Compatibility of resulting system of algebraic equations (which is linear in coefficients of $\tau_{i+1}$) imposes restrictions on the free parameters of $\tau_{i}$. For sequences related to the even bi-spectral family, there exists a Wronskian representation (see Section 13).
We recall that the main sequences exhaust all possible polynomial solutions of (10) with no common/multiple roots for the cases $\Lambda=1$ and $\Lambda=2$. Solutions of (69) that do not belong to the main sequences do not satisfy the bi-linear equation (10); instead, they satisfy its generalization.
In more detail: Let us first consider a sequence of the Darboux transforms at the zero energy level, starting from $H_{0}=-\partial_{z}^{2}+u_{0}$, with non-zero $u_{0}\not=0$. At the $n$th step of the chain, we have
$u_{n}=u_{0}-2(\log\tau_{n})^{\prime\prime},\quad\varkappa_{n}=\frac{\tau_{n+1}}{\tau_{n}},\quad\tau_{0}=1,\quad H_{n}\varkappa_{n}=0.$
@@
In fact, this generalization is a particular case of the tri-linear equation for three species of vortices with ${\cal Q}_{i}\in\{-1,1,a\}$, where a single vortex of the third species of strength $a$ is placed at the origin $z=0$ [20, 42]. In general, for multiple species we have the poly-linear equation (see e.g. [32, 40])
$\sum_{i}\Lambda_{i}^{2}\frac{{\cal P}_{i}^{\prime\prime}}{{\cal P}_{i}}+2\sum_{i<j}\Lambda_{i}\Lambda_{j}\frac{{\cal P}_{i}^{\prime}}{{\cal P}_{i}}\frac{{\cal P}_{j}^{\prime}}{{\cal P}_{j}}=0,$ (71)
where sums run over all species, $\Lambda_{i}$ is a strength of the $i$th species, and ${\cal P}_{i}$ is polynomial whose roots correspond to the coordinates of vortices of this species. In the case of configurations related to the even bi-spectral family ${\cal P}_{1}=p$, ${\cal P}_{2}=q$, and ${\cal P}_{3}=z$. In a similar situation, when ${\cal Q}_{i}\in\{-1,\Lambda,\Lambda a\}$, the following generalization of Eq. (10) holds